OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log4x2 = log4(5x + 14)

OpenStudy (mathstudent55):

If log x = log y, then x = y.

OpenStudy (campbell_st):

equate the expressions since the logs are equal and the same base \[x^2 = 5x + 14\] so solve the above expression and remember you can't find the log of a number less than zero

OpenStudy (mathstudent55):

^less than or equal to zero

OpenStudy (anonymous):

\[\log_4x^2 = \log_4(5x + 14) \implies x^2-5x-14=(x-7)(x+2)=0\]\[ \implies x=7~~ or~~ x=-2\]However, the second solution won't work, since -2 is not in the domain of the original problem, so the solution is \[x=7\]

OpenStudy (mathstudent55):

@AnimalAin Why will x = -2 not work?

OpenStudy (anonymous):

Excellent point. I guess I only get half credit on this one.....

OpenStudy (anonymous):

Yeah, it works fine. Brain fade, I guess; after my bedtime.

OpenStudy (mathstudent55):

\(x = -2\) \( \log_{4} x^2 = \log_{4} (-2)^2 = \log_{4} 4 = 1 \) \( \log_{4} (5x + 14) = \log_{4} [5(-2) + 14] = \log_{4} (-10 + 14) = \log_{4} 4 = 1\)